OpenStudy (anonymous):

I need to drink 8 2/4 cups of water and 2 1/5 cups of milk every day. How much fluid do i have to drink?

OpenStudy (anonymous):

step by step please

OpenStudy (dumbcow):

Add the mixed numbers 8 +2 = 10 for the fraction part, get common denominator LCM of 4 and 5 is 20 2/4 + 1/5 = 10/20 + 4/20 = 14/20 = 7/10 answer -> 10 7/10

OpenStudy (anonymous):

ty